# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v1

@@ Appendix E.2 RQ2 LLM Stages of Moral Development at All Temperatures
Model Temp CC-FC DJ-SS C-SR L-SN FA-ES Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.97 1.92 2.00 1.94 2.00 1.97 0.25 2.00 2.00 1.92 1.94 2.00 1.98 0.50 1.97 2.00 1.92 2.00 2.00 1.98 0.75 1.97 2.00 2.00 2.00 2.00 1.99 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Claude-3.7 0.00 2.76 2.85 2.50 2.78 2.94 2.77 0.25 2.83 2.85 2.58 2.83 2.94 2.82 0.50 2.83 2.77 2.58 2.89 2.88 2.81 0.75 2.90 2.85 2.58 2.72 2.94 2.82 1.00 2.86 2.77 2.50 2.83 2.94 2.81 R1-Distill-70B 0.00 2.62 2.85 2.58 2.83 2.94 2.75 0.25 2.62 2.85 2.67 2.67 2.75 2.69 0.50 2.69 2.77 2.75 2.67 2.81 2.73 0.75 2.76 2.85 2.58 2.61 2.81 2.73 1.00 2.72 2.85 2.67 2.67 2.69 2.72 QwQ-32B 0.00 2.00 2.00 2.00 2.00 2.00 2.00 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 1.97 1.92 2.00 2.00 2.00 1.98 1.00 2.00 2.00 2.00 2.00 2.00 2.00 R1-Distill-8B 0.00 2.34 2.77 2.67 2.56 2.56 2.53 0.25 2.41 2.54 2.67 2.50 2.69 2.53 0.50 2.48 2.77 2.42 2.50 2.62 2.54 0.75 2.62 2.69 2.50 2.56 2.62 2.60 1.00 2.48 2.85 2.33 2.50 2.38 2.50 GLM-Z1-9B 0.00 2.62 2.77 2.75 2.78 2.94 2.75 0.25 2.69 2.85 2.67 2.67 2.94 2.75 0.50 2.62 2.85 2.58 2.89 2.94 2.76 0.75 2.76 2.77 2.67 2.83 2.94 2.80 1.00 2.79 2.85 2.83 2.78 3.00 2.84 Non-Reasoning Models GPT-4.1 0.00 2.69 2.69 2.67 2.83 2.88 2.75 0.25 2.66 2.77 2.58 2.72 2.81 2.70 0.50 2.59 2.69 2.58 2.78 2.81 2.68 0.75 2.59 2.62 2.58 2.67 2.75 2.64 1.00 2.59 2.62 2.58 2.67 2.75 2.64 DeepSeek-V3 0.00 2.00 2.00 2.00 2.00 2.00 2.00 0.25 1.97 2.00 2.00 2.00 2.00 1.99 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Qwen2.5-72B 0.00 2.00 2.00 2.00 1.94 2.00 1.99 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Qwen2.5-32B 0.00 2.69 2.92 2.58 2.72 2.81 2.74 0.25 2.55 2.77 2.50 2.67 2.75 2.64 0.50 2.45 2.62 2.75 2.72 2.75 2.63 0.75 2.59 2.77 2.58 2.78 2.88 2.71 1.00 2.55 2.69 2.50 2.78 2.75 2.65 Qwen2.5-7B 0.00 2.00 2.00 2.00 1.94 2.00 1.99 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Baichuan2-7B 0.00 2.62 2.54 2.58 2.61 2.88 2.65 0.25 2.55 2.77 2.33 2.61 2.75 2.60 0.50 2.59 2.77 2.25 2.61 2.69 2.59 0.75 2.62 2.62 2.42 2.44 2.75 2.58 1.00 2.48 2.77 2.50 2.50 2.81 2.59
@@
Appendix E.3 RQ3 Harmful Response Rates under Soft Prompt Injection
@@
Appendix Appendix F Use Of AI Assistants
@@
Model Temp INC OFF IND IR Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.00 1.00 0.96 0.71 0.92 0.25 1.00 1.00 1.00 0.40 0.85 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 1.00 0.44 0.86 1.00 1.00 1.00 1.00 0.48 0.87 Claude-3.7 0.00 1.00 0.60 0.96 0.00 0.64 0.25 1.00 0.56 1.00 0.00 0.64 0.50 1.00 0.60 1.00 0.00 0.65 0.75 1.00 0.48 1.00 0.08 0.64 1.00 1.00 0.56 0.96 0.12 0.66 R1-Distill-70B 0.00 0.64 0.92 1.00 0.20 0.69 0.25 0.72 1.00 0.92 0.24 0.72 0.50 0.52 1.00 1.00 0.16 0.67 0.75 0.44 1.00 0.96 0.36 0.69 1.00 0.48 1.00 1.00 0.16 0.66 QwQ-32B 0.00 0.96 1.00 1.00 0.32 0.82 0.25 1.00 1.00 1.00 0.40 0.85 0.50 0.96 1.00 1.00 0.40 0.84 0.75 1.00 1.00 1.00 0.40 0.85 1.00 1.00 1.00 1.00 0.48 0.87 R1-Distill-8B 0.00 0.48 0.80 0.84 0.08 0.55 0.25 0.32 0.84 0.84 0.16 0.54 0.50 0.52 0.80 0.72 0.12 0.54 0.75 0.24 0.80 0.72 0.08 0.46 1.00 0.16 0.76 0.72 0.08 0.43 GLM-Z1-9B 0.00 0.96 1.00 1.00 0.60 0.89 0.25 0.92 1.00 1.00 0.48 0.85 0.50 0.96 1.00 1.00 0.48 0.86 0.75 0.88 1.00 1.00 0.60 0.87 1.00 1.00 1.00 1.00 0.60 0.90 Non-Reasoning Models GPT-4.1 0.00 1.00 1.00 0.84 0.44 0.86 0.25 1.00 1.00 0.88 0.44 0.87 0.50 0.96 1.00 0.92 0.41 0.86 0.75 1.00 1.00 0.88 0.44 0.86 1.00 1.00 1.00 0.92 0.33 0.85 DeepSeek-V3 0.00 1.00 1.00 1.00 0.20 0.80 0.25 1.00 1.00 0.96 0.42 0.85 0.50 1.00 1.00 1.00 0.12 0.78 0.75 1.00 1.00 1.00 0.20 0.80 1.00 1.00 1.00 1.00 0.12 0.78 Qwen2.5-72B 0.00 0.60 0.32 0.52 0.00 0.36 0.25 0.72 0.36 0.52 0.00 0.40 0.50 0.60 0.44 0.60 0.00 0.41 0.75 0.52 0.52 0.60 0.00 0.41 1.00 0.52 0.60 0.48 0.00 0.40 Qwen2.5-32B 0.00 0.36 0.96 0.52 0.00 0.46 0.25 0.32 0.92 0.52 0.00 0.44 0.50 0.20 0.96 0.56 0.00 0.43 0.75 0.32 0.96 0.44 0.00 0.43 1.00 0.36 0.96 0.48 0.00 0.45 Qwen2.5-7B 0.00 0.08 0.52 0.28 0.00 0.22 0.25 0.08 0.48 0.20 0.04 0.20 0.50 0.12 0.48 0.20 0.04 0.21 0.75 0.08 0.40 0.20 0.00 0.17 1.00 0.08 0.44 0.12 0.04 0.17 Baichuan2-7B 0.00 0.28 0.12 0.08 0.04 0.13 0.25 0.32 0.12 0.28 0.12 0.21 0.50 0.12 0.16 0.16 0.04 0.12 0.75 0.16 0.20 0.24 0.00 0.15 1.00 0.16 0.16 0.12 0.00 0.11

### v2

@@ Appendix H.2 RQ2 LLM Stages of Moral Development at All Temperatures
Model Temp CC-FC DJ-SS C-SR L-SN FA-ES Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.97 1.92 2.00 1.94 2.00 1.97 0.25 2.00 2.00 1.92 1.94 2.00 1.97 0.50 1.97 2.00 1.92 2.00 2.00 1.98 0.75 1.97 2.00 2.00 2.00 2.00 1.99 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Claude-3.7 0.00 2.76 2.85 2.50 2.78 2.94 2.76 0.25 2.83 2.85 2.58 2.83 2.94 2.81 0.50 2.83 2.77 2.58 2.89 2.88 2.79 0.75 2.90 2.85 2.58 2.72 2.94 2.80 1.00 2.86 2.77 2.50 2.83 2.94 2.78 R1-Distill-70B 0.00 2.62 2.85 2.58 2.83 2.94 2.76 0.25 2.62 2.85 2.67 2.67 2.75 2.71 0.50 2.69 2.77 2.75 2.67 2.81 2.74 0.75 2.76 2.85 2.58 2.61 2.81 2.72 1.00 2.72 2.85 2.67 2.67 2.69 2.72 QwQ-32B 0.00 2.00 2.00 2.00 2.00 2.00 2.00 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 1.97 1.92 2.00 2.00 2.00 1.98 1.00 2.00 2.00 2.00 2.00 2.00 2.00 R1-Distill-8B 0.00 2.34 2.77 2.67 2.56 2.56 2.58 0.25 2.41 2.54 2.67 2.50 2.69 2.56 0.50 2.48 2.77 2.42 2.50 2.62 2.56 0.75 2.62 2.69 2.50 2.56 2.62 2.60 1.00 2.48 2.85 2.33 2.50 2.38 2.51 GLM-Z1-9B 0.00 2.62 2.77 2.75 2.78 2.94 2.77 0.25 2.69 2.85 2.67 2.67 2.94 2.76 0.50 2.62 2.85 2.58 2.89 2.94 2.78 0.75 2.76 2.77 2.67 2.83 2.94 2.79 1.00 2.79 2.85 2.83 2.78 3.00 2.85 Qwen3-235B-R 0.00 2.79 2.85 2.58 2.89 2.94 2.81 0.25 2.76 2.85 2.67 2.83 2.88 2.80 0.50 2.86 2.85 2.67 2.89 3.00 2.85 0.75 2.86 2.85 2.75 2.89 2.94 2.86 1.00 2.76 2.85 2.67 2.89 3.00 2.83 Non-Reasoning Models GPT-4.1 0.00 2.69 2.69 2.67 2.83 2.88 2.75 0.25 2.66 2.77 2.58 2.72 2.81 2.71 0.50 2.59 2.69 2.58 2.78 2.81 2.69 0.75 2.59 2.62 2.58 2.67 2.75 2.64 1.00 2.59 2.62 2.58 2.67 2.75 2.64 DeepSeek-V3 0.00 2.00 2.00 2.00 2.00 2.00 2.00 0.25 1.97 2.00 2.00 2.00 2.00 1.99 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Qwen2.5-72B 0.00 2.00 2.00 2.00 1.94 2.00 1.99 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Qwen2.5-32B 0.00 2.69 2.92 2.58 2.72 2.81 2.75 0.25 2.55 2.77 2.50 2.67 2.75 2.65 0.50 2.45 2.62 2.75 2.72 2.75 2.66 0.75 2.59 2.77 2.58 2.78 2.88 2.72 1.00 2.55 2.69 2.50 2.78 2.75 2.65 Qwen2.5-7B 0.00 2.00 2.00 2.00 1.94 2.00 1.99 0.25 2.00 2.00 2.00 2.00 2.00 2.00 0.50 2.00 2.00 2.00 2.00 2.00 2.00 0.75 2.00 2.00 2.00 2.00 2.00 2.00 1.00 2.00 2.00 2.00 2.00 2.00 2.00 Baichuan2-7B 0.00 2.62 2.54 2.58 2.61 2.88 2.65 0.25 2.55 2.77 2.33 2.61 2.75 2.60 0.50 2.59 2.77 2.25 2.61 2.69 2.58 0.75 2.62 2.62 2.42 2.44 2.75 2.57 1.00 2.48 2.77 2.50 2.50 2.81 2.61 Qwen3-235B-C 0.00 2.62 3.00 2.50 2.61 2.69 2.68 0.25 2.62 3.00 2.50 2.61 2.69 2.68 0.50 2.62 3.00 2.50 2.61 2.69 2.68 0.75 2.62 3.00 2.50 2.61 2.69 2.68 1.00 2.62 3.00 2.50 2.61 2.69 2.68
Table 16: MSS of 14 LLMs Across 5 Moral Dimensions (Chinese, temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (88 dilemmas)
Model Temp CC-FC DJ-SS C-SR L-SN FA-ES Overall Reasoning-Oriented Models DeepSeek-R1 0.00 3.00 2.85 2.92 2.83 3.00 2.92 0.25 2.93 2.77 2.75 2.89 3.00 2.87 0.50 2.93 2.85 3.00 2.83 3.00 2.92 0.75 2.93 2.85 2.92 2.83 3.00 2.91 1.00 2.97 2.85 2.92 2.83 3.00 2.91 Claude-3.7 0.00 3.00 2.85 2.83 2.72 2.92 2.86 0.25 2.97 2.77 2.92 2.67 3.00 2.86 0.50 2.97 2.77 2.92 2.83 2.93 2.88 0.75 2.93 2.77 2.92 2.78 2.87 2.85 1.00 2.83 2.92 2.83 2.83 2.80 2.84 R1-Distill-70B 0.00 2.69 2.77 2.58 2.50 2.62 2.63 0.25 2.48 2.85 2.58 2.56 2.62 2.62 0.50 2.66 2.85 2.50 2.56 2.69 2.65 0.75 2.59 2.85 2.67 2.44 2.75 2.66 1.00 2.66 2.85 2.58 2.61 2.75 2.69 QwQ-32B 0.00 2.93 2.85 2.92 2.72 3.00 2.88 0.25 2.83 2.77 2.75 2.83 3.00 2.84 0.50 2.86 2.77 2.75 2.78 2.94 2.82 0.75 2.83 2.62 2.92 2.89 3.00 2.85 1.00 2.86 2.77 2.75 2.72 2.94 2.81 R1-Distill-8B 0.00 2.45 2.46 2.42 2.39 2.69 2.48 0.25 2.45 2.54 2.42 2.44 2.69 2.51 0.50 2.59 2.46 2.42 2.56 2.75 2.55 0.75 2.41 2.62 2.42 2.44 2.69 2.52 1.00 2.34 2.31 2.50 2.44 2.69 2.46 GLM-Z1-9B 0.00 2.90 2.77 2.83 2.83 3.00 2.87 0.25 2.62 3.00 2.50 2.61 2.69 2.68 0.50 2.62 3.00 2.50 2.61 2.69 2.68 0.75 2.62 3.00 2.50 2.61 2.69 2.68 1.00 2.93 2.85 2.92 2.89 2.94 2.90 Qwen3-235B-R 0.00 2.83 2.85 3.00 2.83 3.00 2.90 0.25 3.00 2.77 2.92 2.83 3.00 2.90 0.50 2.93 2.85 2.92 2.89 2.88 2.89 0.75 2.90 2.92 2.92 2.83 2.94 2.90 1.00 2.90 2.85 2.92 2.94 3.00 2.92 Non-Reasoning Models GPT-4.1 0.00 2.59 3.00 2.50 2.50 2.62 2.64 0.25 2.59 3.00 2.50 2.50 2.62 2.64 0.50 2.59 3.00 2.50 2.50 2.62 2.64 0.75 2.59 2.92 2.58 2.50 2.60 2.64 1.00 2.59 3.00 2.50 2.50 2.62 2.64 DeepSeek-V3 0.00 2.97 2.92 3.00 2.78 3.00 2.93 0.25 2.83 2.85 2.92 2.72 2.94 2.85 0.50 2.93 3.00 3.00 2.78 2.94 2.93 0.75 2.93 2.92 2.83 2.83 3.00 2.90 1.00 2.90 2.85 3.00 2.78 2.94 2.89 Qwen2.5-72B 0.00 2.86 2.92 2.75 2.83 2.81 2.84 0.25 2.83 2.85 2.67 2.83 2.81 2.80 0.50 2.83 3.00 2.67 2.72 2.94 2.83 0.75 2.83 2.85 2.58 2.83 2.88 2.79 1.00 2.90 3.00 2.67 2.78 2.94 2.86 Qwen2.5-32B 0.00 2.72 2.77 2.67 2.61 2.81 2.72 0.25 2.76 2.92 2.67 2.56 2.75 2.73 0.50 2.66 2.85 2.75 2.61 2.75 2.72 0.75 2.72 2.92 2.67 2.72 2.81 2.77 1.00 2.66 2.77 2.67 2.56 2.94 2.72 Qwen2.5-7B 0.00 2.93 2.92 2.92 2.78 2.88 2.88 0.25 2.79 2.92 2.83 2.78 2.81 2.83 0.50 2.86 2.85 2.83 2.78 2.94 2.85 0.75 2.83 2.92 2.83 2.83 2.88 2.86 1.00 2.83 3.00 2.83 2.72 2.88 2.85 Baichuan2-7B 0.00 2.73 2.82 2.62 2.74 2.73 2.73 0.25 2.73 2.83 2.65 2.65 2.74 2.72 0.50 2.74 2.83 2.64 2.64 2.74 2.72 0.75 2.75 2.84 2.67 2.64 2.75 2.73 1.00 2.75 2.85 2.68 2.68 2.75 2.74 Qwen3-235B-C 0.00 2.62 3.00 2.50 2.61 2.69 2.68 0.25 2.62 3.00 2.50 2.61 2.69 2.68 0.50 2.62 3.00 2.50 2.61 2.69 2.68 0.75 2.62 3.00 2.50 2.61 2.69 2.68 1.00 2.62 3.00 2.50 2.61 2.69 2.68
@@
Appendix H.3 RQ3 Harmful Response Rates under Soft Prompt Injection
@@
Appendix Appendix I Use Of AI Assistants
@@
Model Temp INC OFF IND IR Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.00 1.00 0.96 0.71 0.92 0.25 1.00 1.00 1.00 0.40 0.85 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 1.00 0.44 0.86 1.00 1.00 1.00 1.00 0.48 0.87 Claude-3.7 0.00 1.00 0.60 0.96 0.00 0.64 0.25 1.00 0.56 1.00 0.00 0.64 0.50 1.00 0.60 1.00 0.00 0.65 0.75 1.00 0.48 1.00 0.08 0.64 1.00 1.00 0.56 0.96 0.12 0.66 R1-Distill-70B 0.00 0.64 0.92 1.00 0.20 0.69 0.25 0.72 1.00 0.92 0.24 0.72 0.50 0.52 1.00 1.00 0.16 0.67 0.75 0.44 1.00 0.96 0.36 0.69 1.00 0.48 1.00 1.00 0.16 0.66 QwQ-32B 0.00 0.96 1.00 1.00 0.32 0.82 0.25 1.00 1.00 1.00 0.40 0.85 0.50 0.96 1.00 1.00 0.40 0.84 0.75 1.00 1.00 1.00 0.40 0.85 1.00 1.00 1.00 1.00 0.48 0.87 R1-Distill-8B 0.00 0.48 0.80 0.84 0.08 0.55 0.25 0.32 0.84 0.84 0.16 0.54 0.50 0.52 0.80 0.72 0.12 0.54 0.75 0.24 0.80 0.72 0.08 0.46 1.00 0.16 0.76 0.72 0.08 0.43 GLM-Z1-9B 0.00 0.96 1.00 1.00 0.60 0.89 0.25 0.92 1.00 1.00 0.48 0.85 0.50 0.96 1.00 1.00 0.48 0.86 0.75 0.88 1.00 1.00 0.60 0.87 1.00 1.00 1.00 1.00 0.60 0.90 Qwen3-235B-R 0.00 1.00 1.00 1.00 0.56 0.89 0.25 1.00 1.00 1.00 0.48 0.87 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 0.96 0.60 0.89 1.00 1.00 1.00 1.00 0.44 0.86 Non-Reasoning Models GPT-4.1 0.00 1.00 1.00 0.84 0.44 0.86 0.25 1.00 1.00 0.88 0.44 0.87 0.50 0.96 1.00 0.92 0.41 0.86 0.75 1.00 1.00 0.88 0.44 0.86 1.00 1.00 1.00 0.92 0.33 0.85 DeepSeek-V3 0.00 1.00 1.00 1.00 0.20 0.80 0.25 1.00 1.00 0.96 0.42 0.85 0.50 1.00 1.00 1.00 0.12 0.78 0.75 1.00 1.00 1.00 0.20 0.80 1.00 1.00 1.00 1.00 0.12 0.78 Qwen2.5-72B 0.00 0.60 0.32 0.52 0.00 0.36 0.25 0.72 0.36 0.52 0.00 0.40 0.50 0.60 0.44 0.60 0.00 0.41 0.75 0.52 0.52 0.60 0.00 0.41 1.00 0.52 0.60 0.48 0.00 0.40 Qwen2.5-32B 0.00 0.36 0.96 0.52 0.00 0.46 0.25 0.32 0.92 0.52 0.00 0.44 0.50 0.20 0.96 0.56 0.00 0.43 0.75 0.32 0.96 0.44 0.00 0.43 1.00 0.36 0.96 0.48 0.00 0.45 Qwen2.5-7B 0.00 0.08 0.52 0.28 0.00 0.22 0.25 0.08 0.48 0.20 0.04 0.20 0.50 0.12 0.48 0.20 0.04 0.21 0.75 0.08 0.40 0.20 0.00 0.17 1.00 0.08 0.44 0.12 0.04 0.17 Baichuan2-7B 0.00 0.28 0.12 0.08 0.04 0.13 0.25 0.32 0.12 0.28 0.12 0.21 0.50 0.12 0.16 0.16 0.04 0.12 0.75 0.16 0.20 0.24 0.00 0.15 1.00 0.16 0.16 0.12 0.00 0.11 Qwen3-235B-C 0.00 0.84 1.00 0.88 0.00 0.68 0.25 0.84 1.00 0.92 0.00 0.69 0.50 0.96 1.00 0.84 0.08 0.72 0.75 0.84 1.00 0.84 0.00 0.67 1.00 0.96 1.00 0.80 0.00 0.69